OpenStudy (anonymous):

Solve. Please/thanks! -2/z + 1/4z = 7/z+10 A) z=10/3 B) z=-2 C) z=2 D) z=-5/4

OpenStudy (anonymous):

What have you done so far regarding this question?

OpenStudy (dwelsher):

-2/z+1/4z=7/z=10 multiply both sides by 4z :::::-8+1=28+40z 40z=-35 z=-7/8

OpenStudy (dwelsher):

-2/z+1/4z=7/z+10 on the problem above this is the correct original problem ---sorry try to find a common denominator- which is 4z Is the correct answer???

OpenStudy (anonymous):

\[\large -\frac{2}{z}+\frac{1}{4z}=\frac{7}{z+10}\] \[\large \frac{-8z+z}{4z^2}=\frac{7}{z+10}\] \[\large -\frac{7}{4z}=\frac{7}{z+10}\] \[\large -7(z+10)=28z\] \[\large -7z-70=28z\] \[\large 35z=-70\] \[\large z=-2\]